# Video Signature: In-generation Watermarking for Latent Video Diffusion Models

### v1

@@ Abstract
The rapid development of Artificial Intelligence Generated Content (AIGC) has led to significant progress in video generation but also raises serious concerns about intellectual property protection and reliable content tracing. Watermarking is a widely adopted solution to this issue, but existing methods for video generation mainly follow a post-generation paradigm, which introduces additional computational overhead and often fails to effectively balance the trade-off between video quality and watermark extraction. To address these issues, we propose Video Signature (VidSig), an in-generation watermarking method for latent video diffusion models, which enables implicit and adaptive watermark integration during generation. Specifically, we achieve this by partially fine-tuning the latent decoder, where Perturbation-Aware Suppression (PAS) pre-identifies and freezes perceptually sensitive layers to preserve visual quality. Beyond spatial fidelity, we further enhance temporal consistency by introducing a lightweight Temporal Alignment module that guides the decoder to generate coherent frame sequences during fine-tuning. Experimental results show that VidSig achieves the best overall performance in watermark extraction, visual quality, and generation efficiency. It also demonstrates strong robustness against both spatial and temporal tampering, highlighting its practicality in real-world scenarios.
@@ 4.4 Robustness
Model Method Benign FD FS FI FIG FA (n=3) FA (n=5) FA (n=7) FA (n=9) Avg
@@ 5 Conclusion and Limitation
Figure 7: Qualitative comparison of the original and watermarked videos. The first row is the original video, the second row is the video generated by VidSig, and the third row refers to the pixel-wise difference (×$\times 10$).
@@ A.1 Dataset
As we mentioned in Section 4.1, we select 50 prompts from the test set of VBench [Huang et al., 2024b], covering five categories: Animal, Human, Plant, Scenery, and Vehicles (10 prompts each). We list these prompts in Table 5. For each prompt, we generate four videos using fixed random seeds (42, 114514, 3407, $6$) to ensure reproducibility. For the I2V task, we first generate an image using Stable Diffusion 2.1 [Rombach et al., 2022], and subsequently generate four videos conditioned on each generated image using the same set of random seeds. Each generated video contains 16 frames with a resolution $512\times 512$.
@@ A.4 Temporal Tampering implementation
In this paper, to test the resilience of VidSig, we implement several temporal attacks. We provide detailed descriptions of these attacks here. Give a video $[f_{1},f_{2},f_{3},\ldots,f_{N}]$ consists of $N$ frames, we define the following temporal attacks.
@@
Randomly select a frame and delete it. Let $i$ denote the index of the selected frame, the tampered video is then denoted as $[f_{1},f_{2},f_{i-1},f_{i+1},\ldots,f_{N}]$
@@
We randomly select two distinct frames, indexed by $i$ and $j$, and swap their positions in the video, resulting in a perturbed sequence $[f_{1},\ldots,f_{i-1},f_{j},\ldots,f_{j-1},f_{i},\ldots,f_{N}]$.
@@
We randomly select two positions indexed by $i$ and $j$, and insert the $i^{\text{th}}$ frame at position $j$, resulting in a new sequence with duplicated content and increased temporal length $[f_{1},f_{2},f_{i},\ldots f_{j-1},f_{i},f_{j}\ldots,f_{N+1}]$
@@
Randomly select a position indexed by $i$, and insert a standard Gaussian noise at position $i$, denoted as $f_{i}^{\prime}$, resulting in a new sequence with increased temporal length $[f_{1},f_{2},\ldots,f_{i}^{\prime},f_{i}\ldots,f_{N+1}]$
@@
Given a sequence length $n$, we randomly select a position indexed by $i$, conditioned on $i+n\leq N$. We then compute the average of the $n$ subsequent frames, i.e., $\bar{f}=\frac{1}{n}\sum_{j=0}^{n-1}f_{i+j}$, and replace the $i^{\text{th}}$ frame with $\bar{f}$, resulting the tampered video $[f_{1},f_{2},\ldots,f_{i-1},\bar{f},\ldots,f_{N-n+1}]$
@@ B.1 Watermark Detection
$M(\mathbf{m},\mathbf{m}^{\prime})\geq\tau,\quad\text{where }\tau\in\{0,\dots,k\}.$ (9)
@@ B.2 Video Quality
[…]
@@
Given a video $[f_{0},f_{1},f_{2},\ldots,f_{2n}]$, the odd-numbered frames $[f_{1},f_{3},\ldots,f_{2n-1}]$ are dropped, and the remaining even-numbered frames $[f_{0},f_{2},\ldots,f_{2n}]$ are used to interpolate [Li et al., 2023] intermediate frames $[\hat{f}_{1},\hat{f}_{3},\ldots,\hat{f}_{2n-1}]$. The Mean Absolute Error (MAE) between interpolated and original dropped frames is computed and normalized to $[0,1]$, with a larger value indicating better smoothness.

### v2

@@ Abstract
The rapid development of Artificial Intelligence Generated Content (AIGC) has led to significant progress in video generation but also raises serious concerns about intellectual property protection and reliable content tracing. Watermarking is a widely adopted solution to this issue, but existing methods for video generation mainly follow a post-generation paradigm, which introduces additional computational overhead and often fails to effectively balance the trade-off between video quality and watermark extraction. To address these issues, we propose Video Signature (VidSig), an in-generation watermarking method for latent video diffusion models, which enables implicit and adaptive watermark integration during generation. Specifically, we achieve this by partially fine-tuning the latent decoder, where Perturbation-Aware Suppression (PAS) pre-identifies and freezes perceptually sensitive layers to preserve visual quality. Beyond spatial fidelity, we further enhance temporal consistency by introducing a lightweight Temporal Alignment module that guides the decoder to generate coherent frame sequences during fine-tuning. Experimental results show that VidSig achieves the best overall performance in watermark extraction, visual quality, and generation efficiency. It also demonstrates strong robustness against both spatial and temporal tampering, highlighting its practicality in real-world scenarios. Our code is available at here
@@ 4.4 Robustness
Model Method Benign FD FS FI FIG
FA
(n=3)
FA
(n=5)
FA
(n=7)
FA
(n=9)
Avg
@@ 5 Conclusion and Limitation
Figure 7: Qualitative comparison of the original and watermarked videos. The first row is the original video, the second row is the video generated by VidSig, and the third row refers to the pixel-wise difference ($\times 10$).
@@ A.1 Dataset
As we mentioned in Section 4.1, we select 50 prompts from the test set of VBench [Huang et al., 2024b], covering five categories: Animal, Human, Plant, Scenery, and Vehicles (10 prompts each). We list these prompts in Table 5. For each prompt, we generate four videos using fixed random seeds (42, 114514, 3407, 6666) to ensure reproducibility. For the I2V task, we first generate an image using Stable Diffusion 2.1 [Rombach et al., 2022], and subsequently generate four videos conditioned on each generated image using the same set of random seeds. Each generated video contains 16 frames with a resolution $512\times 512$.
@@ A.4 Temporal Tampering implementation
In this paper, to test the resilience of VidSig, we implement several temporal attacks. We provide detailed descriptions of these attacks here. Give a video $f_{1},f_{2},f_{3},\ldots,f_{N}]$ consists of NN frames, we define the following temporal attacks.
@@
Randomly select a frame and delete it. Let $i$ denote the index of the selected frame, the tampered video is then denoted as $f_{1},f_{2},f_{i-1},f_{i+1},\ldots,f_{N}]$
@@
We randomly select two distinct frames, indexed by $i$ and $j$, and swap their positions in the video, resulting in a perturbed sequence $f_{1},\ldots,f_{i-1},f_{j},\ldots,f_{j-1},f_{i},\ldots,f_{N}]$.
@@
We randomly select two positions indexed by $i$ and $j$, and insert the $i^{\text{th}}$ frame at position $j$, resulting in a new sequence with duplicated content and increased temporal length $f_{1},f_{2},f_{i},\ldots f_{j-1},f_{i},f_{j}\ldots,f_{N+1}]$
@@
Randomly select a position indexed by $i$, and insert a standard Gaussian noise at position $i$, denoted as $f_{i}^{\prime}$, resulting in a new sequence with increased temporal length $f_{1},f_{2},\ldots,f_{i}^{\prime},f_{i}\ldots,f_{N+1}]$
@@
Given a sequence length $n$, we randomly select a position indexed by $i$, conditioned on $i+n\leq N$. We then compute the average of the $n$ subsequent frames, i.e., $\bar{f}=\frac{1}{n}\sum_{j=0}^{n-1}f_{i+j}$, and replace the $i^{\text{th}}$ frame with $\bar{f}$, resulting the tampered video $f_{1},f_{2},\ldots,f_{i-1},\bar{f},\ldots,f_{N-n+1}]$
@@ B.1 Watermark Detection
[…]
@@ B.2 Video Quality
where $d_{i}$ is the normalized DINO image feature of the $i^{th}$ frame, and $\langle\cdot,\cdot\rangle$ denotes cosine similarity via dot product.
@@
Given a video $f_{0},f_{1},f_{2},\ldots,f_{2n}]$, the odd-numbered frames $f_{1},f_{3},\ldots,f_{2n-1}]$ are dropped, and the remaining even-numbered frames $f_{0},f_{2},\ldots,f_{2n}]$ are used to interpolate [Li et al., 2023] intermediate frames $\hat{f}_{1},\hat{f}_{3},\ldots,\hat{f}_{2n-1}]$. The Mean Absolute Error (MAE) between interpolated and original dropped frames is computed and normalized to $[0,1]$, with a larger value indicating better smoothness.